OpenStudy (anonymous):

Solve for x: the quantity of x plus 16 all over 3 = 3x

OpenStudy (anonymous):

\[(x+16) \over 3\]

OpenStudy (anonymous):

=3x

OpenStudy (akashdeepdeb):

\[\frac{x+16}{3} = 3x\] x + 16 = 3x * 3 x+16 = 9x 16 = 9x-x 16 = 8x x = 2 Understood? :)

OpenStudy (anonymous):

yes thank you!

OpenStudy (akashdeepdeb):

:)